OpenStudy (anonymous):

I) 4x - (9 - 3x) = 8x-1. Could someone please explain how to do this question to me? Thanks in advance.

jimthompson5910 (jim_thompson5910):

4x - (9 - 3x) = 8x-1 4x - 9 + 3x = 8x-1 ... Distribute 7x - 9 = 8x-1 ... Combine like terms What's next?

OpenStudy (anonymous):

7x with 8x and 9 with 1?

jimthompson5910 (jim_thompson5910):

yes you need to get that 8x over, how?

OpenStudy (anonymous):

add with 7 makes 13x

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

subtract 8x from both sides